OpenStudy (anonymous):

In any triangle abc the value of a(b^2+c^2)cos A+b (c^2+a^2)cos B+c(a^2+b^2)cos C is

OpenStudy (anonymous):

c^2 = a^2 + b^2 - 2ab cos(C) b^2 = a^2 + c^2 - 2ac cos(B) a^2 = b^2 + c^2 - 2bc cos(A)

OpenStudy (anonymous):

cos A = (b² +c² -a²)/2bc cos B = (a² +c² -b²)/2ca cos C = (a² +b² -c²)/2ab

OpenStudy (anonymous):

c^2+2abcosc=a^2+b^2

OpenStudy (anonymous):

b^2+2ac cos(B)=a^2 + c^2

OpenStudy (anonymous):

a^2 +2bc cos(A) = b^2 + c^2

OpenStudy (anonymous):

a(a^2+2bc cos(a))cos a is it correct? plug b^2+c^2 value

ganeshie8 (ganeshie8):

i am getting \(3abc\) as final answer after lot of algebra, there must be some simpler way >.<

OpenStudy (anonymous):

yes @ganeshie8 3abc is the correct answer

OpenStudy (anonymous):

u got it xD;)

ganeshie8 (ganeshie8):

yes but it took me two pages :o

ganeshie8 (ganeshie8):

I have used these : cos A = (b² +c² -a²)/2bc cos B = (a² +c² -b²)/2ca cos C = (a² +b² -c²)/2ab

OpenStudy (anonymous):

oh show the computation

OpenStudy (anonymous):

heyy i got an idea whether it is right or not i dont know

OpenStudy (anonymous):

a/sin A = b/sin B = c/sin C

OpenStudy (anonymous):

oh sorry its sine rule

ganeshie8 (ganeshie8):

\[ \begin{array} \\ &= a(b^2+c^2)\cos A+b (c^2+a^2)\cos B+c(a^2+b^2)\cos C \\ &= a(b^2+c^2)\dfrac{b^2+c^2-a^2}{2bc}+b (c^2+a^2)\dfrac{c^2+a^2-b^2}{2ca}+c(a^2+b^2)\dfrac{a^2+b^2-c^2}{2ab} \\ &= a^2(b^2+c^2)\dfrac{b^2+c^2-a^2}{2abc}+b^2 (c^2+a^2)\dfrac{c^2+a^2-b^2}{2abc}+c^2(a^2+b^2)\dfrac{a^2+b^2-c^2}{2abc} \\ \end{array} \]

OpenStudy (anonymous):

i thought that but i wont get answer if its lengthy

OpenStudy (anonymous):

u great done this

ganeshie8 (ganeshie8):

\[ \begin{array} \\ &= \dfrac{1}{2abc} \left( a^2(b^2+c^2)(b^2+c^2-a^2)\color{red}{+}b^2 (c^2+a^2)(c^2+a^2-b^2)\color{red}{+}c^2(a^2+b^2)(a^2+b^2-c^2) \right) \\ \end{array} \]

ganeshie8 (ganeshie8):

im sure there is a better way to do this -.-

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(a/cosA)=(b/cosB)=(c/cosC)

OpenStudy (anonymous):

cosA/a + cosB/b +cosC/c =(a^2 +b^2 - c^2)/2ab

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

\[a^3 \cos A+b^3 \cos B+c^3 \cos C\]

OpenStudy (anonymous):